OpenStudy (anonymous):

plz help me i need help on an 8th grade class

OpenStudy (anonymous):

What is the approximate distance between points A and B?

OpenStudy (anonymous):

3.61 9.22 10.35 12.62

OpenStudy (anonymous):

@aaronq @abb0t @abb0t

OpenStudy (anonymous):

do you know the distance formula?

OpenStudy (joannablackwelder):

Do you know the pythagorean theorem?

OpenStudy (joannablackwelder):

a^2 + b^2 = c^2 for right triangles

OpenStudy (joannablackwelder):

@ShawnSpencer ?

OpenStudy (anonymous):

ok

OpenStudy (joannablackwelder):

So, if we draw a right triangle along the x and y lines, how long are the sides of the right triangle?

OpenStudy (anonymous):

would it be 6

OpenStudy (joannablackwelder):

One side is 6. How about the other side?

OpenStudy (anonymous):

6 1/2

OpenStudy (joannablackwelder):

The side along the line y=-4 is 6. How about the side along the line x=4?

OpenStudy (anonymous):

is the other side

OpenStudy (anonymous):

idk

OpenStudy (joannablackwelder):

From -4 to 3 is a length of 7

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

So, a=6, b=7

OpenStudy (joannablackwelder):

:-)

OpenStudy (joannablackwelder):

So, can you solve for c?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 85

OpenStudy (joannablackwelder):

Close, square root of 85

OpenStudy (joannablackwelder):

Because 85 = c^2

OpenStudy (anonymous):

ok so 7225 ???

OpenStudy (joannablackwelder):

No, 

OpenStudy (anonymous):

ok so how do i do that

OpenStudy (joannablackwelder):

Do you have a calculator?

OpenStudy (anonymous):

ya

OpenStudy (joannablackwelder):

Does it have a square root button?

OpenStudy (joannablackwelder):

Or a y^x button?

OpenStudy (anonymous):

no sorry its like one of those simple ones

OpenStudy (joannablackwelder):

You should have a calc on your computer that you can put into scientific mode. You may need to get a scientific calc for school sometime.

OpenStudy (joannablackwelder):

Open the calc on your computer, go to view and scientific